OpenStudy (anonymous):

Molly went to the beach and collected a bag of seashells. She kept half of her collection. She gave the other half to Beth and Jack. Jack received three times as many seashells as Beth. If Beth received 8 seashells, how many seashells did Molly collect at the beach?

OpenStudy (anonymous):

I need help please

OpenStudy (anonymous):

Beth+Jack = 1/2 * Molly Jack = 3 * Beth Beth = 8 Substituting 3rd equation into second Jack = 3 * 8 = 24 Substituting now into first equation 8 + 24 = 1/2 * Molly ... 32 = 1/2 Molly Multiply both sides by 2 64 = Molly So yes, the answer is 64.